OpenStudy (jennychan12):

integral from 0 to 1 of [e^x(cos(e^x))]dx ?

OpenStudy (jennychan12):

\[\int\limits_{0}^{1} [e^x \cos (e^x)]dx\]

OpenStudy (kinggeorge):

Do a u-substitution with \(u=\sin(e^x)\), and your solution should pop right out.

hartnn (hartnn):

did you mean u=e^x ?

OpenStudy (jennychan12):

but if you do that, then du = e^xcos(e^x) and there's no cos (e^x) in the question.

hartnn (hartnn):

try u=e^x

OpenStudy (jennychan12):

that'd just be ucosu

OpenStudy (kinggeorge):

\(u=\sin(e^x)\implies du=e^x\cos(e^x)dx\). So your integral becomes\[\int du=u\]

OpenStudy (kinggeorge):

Substitute back for \(u\), and you get \[\int_0^1 e^x\cos(e^x)dx=\sin(e^x)|_0^1\]

hartnn (hartnn):

u =e^x du = e^x dx \(\int cos udu\) ohh..now i could say that u =sin e^x is a better substitution..

OpenStudy (jennychan12):

oh wait. my bad. i thought u said u = sin u ok i see now

OpenStudy (kinggeorge):

Yup, I like to call my method "guessing the solution and proving you're right"

OpenStudy (kinggeorge):

Although if you really had no clue of the solution, \(u=e^x\) would be a fine substitution. You would just have to integrate by parts.

hartnn (hartnn):

whenever i see a non-standard angle with sin/cos/.. i'll call that as 'bad angle' and put u= bad angle.... example : sin x^2 , cos log x ....

hartnn (hartnn):

how do we need integration by parts ?? O.o u= e^x du=e^xdx

OpenStudy (kinggeorge):

Oh. Right. You don't. Ignore my ramblings.